very lost help me answer these

4. Suppose that a certain country has an MPC of 0.9 and a real GDP of $400 billion.

If its investment spending decreases by $4 billion, what will be its new level of real GDP?

5. The data in columns 1 and 2 in the table below are for a private closed economy.

GDP

A.E. Private Closed Economy

Exports

Imports

Net Exports

A.E. Private Open Economy

200

240

20

30

250

280

20

30

300

320

20

30

350

360

20

30

400

400

20

30

450

440

20

30

500

480

20

30

550

520

20

30

a. Use columns 1 and 2 to determine the equilibrium GDP for this hypothetical economy.

b. Now open up this economy to international trade by including the export and import figures of columns 3 and 4. Fill in columns 5 and 6 and determine the equilibrium GDP for the open economy. What is the change in equilibrium GDP caused by the addition of net exports?

c. What is the multiplier in this example?

7. Refer to columns 1 and 6 in the table for problem 5. Incorporate government into the table by assuming that it plans to tax and spend $20 billion at each possible level of GDP. Also assume that the tax is a personal tax and that government spending does not induce a shift in the private aggregate expenditures schedule.

What is the change in equilibrium GDP caused by the addition of government?
Answer #1

Answer to Question No. 4

The multiplier K = 1/(1-MPC)

so K = 1/(1-0.69) = 1/0.1 = 10 times

K is also expressed as K = Change in income/Change in investment

so 10 = Change in income/4

Change in income = $40 billion

Answer to Question No. 5

GDP A.E. Private Closed Economy Exports Imports Net Exports A.E. Private Open Economy Remarks MPC Multiplier = 1/(1-MPC)
200 240 20 30 -10 230
250 280 20 30 -10 270 0.8 5
300 320 20 30 -10 310 0.8 5
350 360 20 30 -10 350 Equilibrium is reached with AD(Open) = GDP = 350 0.8 5
400 400 20 30 -10 390 Equilibrium is reached with AD(Closed) = GDP=400 0.8 5
450 440 20 30 -10 430 0.8 5
500 480 20 30 -10 470 0.8 5
550 520 20 30 -10 510 0.8 5

Change in GDP by addition of Net Exports is -$50 billion. The new equilibrium GDP has fallen from $400 to $350 billion.

Formulas Used:

AD(closed) = GDP for closed equilibrium, AD(open) = GDP for open equilibrium

MPC = 1/(1-MPC), Net Exports = Exports-Imports